State in words, Newton’s law of Gravitation.

By considering the centripetal force which acts on a planet in a circular orbit, show that T2  is directly proportional to R3 , where T is the time taken for one orbit around the Sun and R is the radius of the orbit.


1
Expert's answer
2021-02-24T12:50:43-0500

Answer

Newton's law of gravitation, statement that any particle of matter in the universe attracts any other with a force varying directly as the product of the masses and inversely as the square of the distance between them as

"F=\frac{Gmm'}{r^2}"

According to kepler law

"\frac{dA}{dt}=\frac{\pi ab}{T}=\frac{L}{2m}"

"T=\frac{2\pi abm}{L}"

"T=\frac{2\pi a^{\frac{3}{2}}}{\sqrt{Gm}}"

This implies

"T^2\propto a^3"